A manufacturer uses a new production method to produce steel rods. A random sample of 17 steel rods resulted in lengths with a standard deviation of 4.7 cm. At the 0.10 significance level, test the claim that the new production method has lengths with a standard deviation different from 3.5 cm, which was the standard deviation for the old method



In one town, monthly incomes men with college degrees are found to have a standard deviation of $650. Use a 0.01 significance level to test the claim that for men without college degrees in that town, incomes have a higher standard deviation. A random sample of 22 men without college degrees resulted in incomes with a standard deviation of $913.
Answer #1

1)

given that m=17, S: 4.7 ) let los be df = n-1=16 3.5 t Hoo Ho # 3.5 m-ns test Statistic: gita x= [17-1) [4.75 3:5) * 28.8522